OpenStudy (morgengarner):

Simplify the expression. (–5g5h6)2(g4h2)4 A. 25g15h14 B. The quantity g raised to the power of 26 multiplied by h raised to the power of 20 end of quantity divided by 25. C. –25g26h20 D. 25g26h20

OpenStudy (nile):

hi

OpenStudy (morgengarner):

hi

jhonyy9 (jhonyy9):

what is there exactly ? so (-5g^5h^6)^2 *(g^4h^2)^4 ?

OpenStudy (morgengarner):

yep ^

jhonyy9 (jhonyy9):

use the rule of exponents (a^x)^y = a^(x*y)

jhonyy9 (jhonyy9):

and (a^x) *(a^y) = a^(x+y)